SANUMARC produces fingerlings for sell. The quantity x (kg) of these fingerlings demanded each week is related to the wholesale unit price p by the equation P = − 0.006x + 180 The weekly total cost incurred by SANUMARC for producing x kgs of fingerlings is C(x) = 0.000002x3 – 0.02x2 + 120x + 60,00 a. Find the marginal cost function C, [5] b. Find the marginal revenue function R’ [5] c. Find the marginal profit function P’ (5) d. Compute P’(2000) and interpret the results.


1
Expert's answer
2021-10-05T09:20:57-0400

a

"c=0.000002x^3-0.02x^2+120x+6000"

marginal cost "=c=\frac{dc}{dx}=0000006x^2-0.04x+120"


b

"P=-0.006x+180"

Revenue=R"=px=-0.006x^2+180x\\"


Marginal Revenue=R'"=\frac{dR}{dx}=-0.012x+180"


c.

Marginal profit"=R'-C'"

"P'=-0.012x+180-0.000006x^2+0.04x-120\\=-0.000006x^2+0.028x+60"


d.

"x=2000\\P'=-0.00000(2000)^2+0.028(2000)+60=92"

This implies that additional profit from selling 2000th unit is 92. Profit is not maximized at this point as marginal profit is not zero